# Strange metal and superconductor in the two-dimensional Yukawa-Sachdev-Ye-Kitaev model

### v1

@@ Abstract
The two-dimensional Yukawa-Sachdev-Ye-Kitaev (YSYK) model provides a universal theory of quantum phase transitions in metals in the presence of quenched random spatial fluctuations in the local position of the quantum critical point. It has a Fermi surface coupled to a scalar field by spatially random Yukawa interactions. We present full numerical solutions of a self-consistent disorder averaged analysis of the YSYK model in both the normal and superconducting states, obtaining electronic spectral functions, frequency-dependent conductivity, and superfluid stiffness. Our results reproduce key aspects of observations in the cuprates as analyzed by Michon et al. (Nat. Comm. 14, 3033 (2023)). We also find a regime of increasing zero temperature superfluid stiffness with decreasing superconducting critical temperature, as is observed in bulk cuprates.
@@
An often quoted model for a strange or bad metal (e.g. [3]) is one in which there is a large density of states of low energy bosonic excitations, usually phonons, and then quasi-elastic scattering of the electrons off the bosons leads to linear-in-$T$ resistivity from the Bose occupation function when $T$ is larger than the typical boson energy. However, studies of the optical conductivity in the strange metal of the cuprates [4] show that the dominant scattering is inelastic, not quasi-elastic, and leads to a non-Drude power-law-in-frequency tail in the optical conductivity. The optical conductivity data has been incisively analyzed recently by Michon et al. [5]: they have shown that while the transport scattering rate (related to the real part of the inverse optical conductivity) exhibits Planckian scaling behavior [1], there are significant logarithmic deviations from scaling in the frequency and temperature dependent effective transport mass (related to the imaginary part of the inverse optical conductivity). Furthermore, the optical conductivity data connects consistently with d.c. measurements of resistivity and thermodynamics.
Our paper presents a self-consistent, disorder-averaged analysis of a two-dimensional Yukawa-Sachdev-Ye-Kitaev (YSYK) model, which has a spatially random Yukawa coupling between fermions, $\psi$, with a Fermi surface and a nearly-critical scalar field, $\phi$. We use methods similar to those which yield the exact solution of the zero-dimensional Sachdev-Ye-Kitaev model. Such a two-dimensional YSYK model has been argued [6, 7, 8] to provide a universal description of quantum phase transitions in metals, associated with the condensation of $\phi$, in the presence of impurity-induced ‘Harris’ disorder [9, 10, 11] with spatial fluctuations in the local position of the quantum critical point. We find results that display all the key characteristics of the optical conductivity and d.c. resistivity described by Michon et al., as shown in Fig. 3.
Moreover, YSYK models also display instabilities of the strange metal to superconductivity [12, 13, 14, 15, 16], with the pairing type dependent upon the particular quantum phase transition being studied. We will examine an instability to spin-singlet pairing in a simplified model which ignores the gap variation around the Fermi surface, and will therefore only apply to the anti-nodal regions of the cuprates. We present the evolution of the electron spectral function and superfluid density for $T<T_{c}$, the superconducting critical temperature. We find that the YSYK model exhibits a number of experimentally observed trends:
@@
The two-dimensional YSYK Model—Several works [21, 22, 23, 24, 25, 26, 27, 28] have argued that clean quantum critical metals cannot serve as a universal model for transport in a strange metal, and that impurity-induced spatial disorder is essential. As a characteristic example, we therefore add Harris disorder to the Hertz-Millis theory of a quantum phase transition in a metal associated with an Ising-nematic order parameter $\phi$ [29]. Such disorder is provided by quenched random terms which preserve the Ising symmetry [23, 24]. Other order parameters, including those at non-zero wavevector, and Fermi-volume changing transitions without broken symmetries [30], also map to essentially the same YSYK model [6]. We define the two-dimensional YSYK model here as the Harris-disordered Hertz-Millis Lagrangian for $\phi$ and fermions $\psi$ with dispersion $\varepsilon({\bm{k}})$ [31, 32, 6, 26, 28]:
$\mathcal{L}_{\rm YSYK}=\psi^{\dagger}_{\sigma}\left[\partial_{\tau}+\varepsilon(i\nabla)-\mu\right]\psi_{\sigma}+v({\bm{r}})\psi^{\dagger}_{\sigma}\psi_{\sigma}\,$
@@
$\mathcal{L}_{\rm YSYK}$ contains the two sources of disorder. One is the potential $v({\bm{r}})$ acting on the fermions:
@@
Two distinct regimes of behavior of $\mathcal{L}_{\rm YSYK}$ have been identified [7, 8]:
@@
(ii) At low $T$ there is a crossover to a regime where bosonic eigenmodes are localized, while the fermionic eigenmodes remain extended [40]. Here, we must treat the disorder in the bosonic sector carefully: by a strong disorder renormalization [11], numerically exactly [7, 8], or map to ‘two-level system’ models [41, 42]. Recent analyses [43, 44] of $v({\bm{r}})$ disorder effects along the lines of Ref. 33 near quantum criticality found singular corrections to the boson propagator, which we view as a precursor to boson localization. This boson localization regime is not described in the present paper.
@@
Figure 3: (a,b) Normal state optical conductivity at the quantum critical coupling for various temperatures indicated in Panel (a). (c,d) Transport scattering rate $1/\tau_{\rm tr}$ and effective mass $m^{*}$ obtained from $\sigma$ using (3); inset of Panel (c) is a scaling plot with $1/\tau_{\rm tr}$ and $\omega$ scaled by $T$, inset of Panel (d) is a scaling plotting showing $\Delta m^{*}/m=m^{*}(\omega)/m-m^{*}(0)/m$ as a function of $\omega/T$. (e,f) Modulus and phase of $\sigma(\omega)$. The effective exponent $\nu^{*}$ is explained in the text. Temperatures ranges are the same for Panels (a)-(d) and for Panels (e) and (f). Our results match the trends of the observations in Figs 3a,3b,1c,2b,1d,2d of Ref. 5
@@
Discussion—The spatially inhomogeneous fermion-boson coupling $g^{\prime}({\bm{r}})$ in the YSYK model in (1) is an alternative to ‘random mass’ $\delta s({\bm{r}})$ disorder in Hertz-Millis theory, in the crossover regime where the bosonic modes remain extended despite the disorder [7], as we discussed below (2). The striking similarities between the results presented here and a variety of properties measured in strange metal superconductors such as the cuprates – including static and dynamic properties across the normal and superconducting states – indicate such disorder plays a significant role in the phenomenology of these systems, making it important to determine the microscopic origin of disorder configurations that strongly affect the local position of the QCP. Finally, the crucial issue of how our results evolve as the bosonic modes localize due to disorder [7] will be addressed in future work.
Acknowledgements—We thank Erez Berg, Christophe Berthod, Andrey Chubukov, Antoine Georges, Peter Lunts, Dmitrii Maslov, Srinivas Raghu, and Dirk van der Marel for valuable discussions. This research was supported by the U.S. National Science Foundation grant No. DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S.S.). The Flatiron Institute is a division of the Simons Foundation. J.S. and D. V. were supported by the German Research Foundation (DFG) through CRC TRR 288 “ElastoQMat,” project A07. I.E. was supported by the University of Wisconsin– Madison

### v2

@@ Abstract
The two-dimensional Yukawa-Sachdev-Ye-Kitaev (2d-YSYK) model provides a universal theory of quantum phase transitions in metals in the presence of quenched random spatial fluctuations in the local position of the quantum critical point. It has a Fermi surface coupled to a scalar field by spatially random Yukawa interactions. We present full numerical solutions of a self-consistent disorder averaged analysis of the 2d-YSYK model in both the normal and superconducting states, obtaining electronic spectral functions, frequency-dependent conductivity, and superfluid stiffness. Our results reproduce key aspects of observations in the cuprates as analyzed by Michon et al. (Nat. Comm. 14, 3033 (2023)). We also find a regime of increasing zero temperature superfluid stiffness with decreasing superconducting critical temperature, as is observed in bulk cuprates.
@@
An often quoted model for a strange or bad metal (e.g. [3, 4]) is one in which there is a large density of states of low energy bosonic excitations, usually phonons, and then quasi-elastic scattering of the electrons off the bosons leads to linear-in-$T$ resistivity from the Bose occupation function when $T$ is larger than the typical boson energy. However, studies of the optical conductivity in the strange metal of the cuprates [5] show that the dominant scattering is inelastic, not quasi-elastic, and leads to a non-Drude power-law-in-frequency tail in the optical conductivity. The optical conductivity data has been incisively analyzed recently by Michon et al. [6]: they have shown that while the transport scattering rate (related to the real part of the inverse optical conductivity) exhibits Planckian scaling behavior [1], there are significant logarithmic deviations from scaling in the frequency and temperature dependent effective transport mass (related to the imaginary part of the inverse optical conductivity). Furthermore, the optical conductivity data connects consistently with d.c. measurements of resistivity and thermodynamics.
Our paper presents a self-consistent, disorder-averaged analysis of a two-dimensional Yukawa-Sachdev-Ye-Kitaev (2d-YSYK) model, which has a spatially random Yukawa coupling between fermions, $\psi$, with a Fermi surface and a nearly-critical scalar field, $\phi$. We use methods similar to those which yield the exact solution of the zero-dimensional Sachdev-Ye-Kitaev model. Such a 2d-YSYK model has been argued [7, 8, 9] to provide a universal description of quantum phase transitions in metals, associated with the condensation of $\phi$, in the presence of impurity-induced ‘Harris’ disorder [10, 11, 12] with spatial fluctuations in the local position of the quantum critical point. We find results that display all the key characteristics of the optical conductivity and d.c. resistivity described by Michon et al., as shown in Fig. 3.
Moreover, YSYK models also display instabilities of the strange metal to superconductivity [13, 14, 15, 16, 17], with the pairing type dependent upon the particular quantum phase transition being studied. We will examine an instability to spin-singlet pairing in a simplified model which ignores the gap variation around the Fermi surface, and will therefore only apply to the anti-nodal regions of the cuprates. We present the evolution of the electron spectral function and superfluid density for $T<T_{c}$, the superconducting critical temperature. We find that the 2d-YSYK model exhibits a number of experimentally observed trends:
@@
The 2d-YSYK Model—Several works [22, 23, 24, 25, 26, 27, 28, 29] have argued that clean quantum critical metals cannot serve as a universal model for transport in a strange metal, and that impurity-induced spatial disorder is essential. As a characteristic example, we therefore add Harris disorder to the Hertz-Millis theory of a quantum phase transition in a metal associated with an Ising-nematic order parameter $\phi$ [30]. Such disorder is provided by quenched random terms which preserve the Ising symmetry [24, 25]. Other order parameters, including those at non-zero wavevector, and Fermi-volume changing transitions without broken symmetries [31], also map to essentially the same 2d-YSYK model [7]. We define the 2d-YSYK model here as the Harris-disordered Hertz-Millis Lagrangian for $\phi$ and fermions $\psi$ with dispersion $\varepsilon({\bm{k}})$ [32, 33, 7, 27, 29, 34]:
$\mathcal{L}_{\rm 2d-YSYK}=\psi^{\dagger}_{\sigma}\left[\partial_{\tau}+\varepsilon(i\nabla)-\mu\right]\psi_{\sigma}+v({\bm{r}})\psi^{\dagger}_{\sigma}\psi_{\sigma}\,$
@@
$\mathcal{L}_{\rm 2d-YSYK}$ contains the two sources of disorder. One is the potential $v({\bm{r}})$ acting on the fermions:
@@
Two distinct regimes of behavior of $\mathcal{L}_{\rm 2d-YSYK}$ have been identified [8, 9]:
@@
(ii) At low $T$ there is a crossover to a regime where bosonic eigenmodes are localized, while the fermionic eigenmodes remain extended [42]. Here, we must treat the disorder in the bosonic sector carefully: by a strong disorder renormalization [12], numerically exactly [8, 9], or map to ‘two-level system’ models [43, 44]. Recent analyses [45, 46] of $v({\bm{r}})$ disorder effects along the lines of Ref. 35 near quantum criticality found singular corrections to the boson propagator, which we view as a precursor to boson localization. This boson localization regime is not described in the present paper.
@@
Figure 3: (a,b) Normal state optical conductivity at the quantum critical coupling for various temperatures indicated in Panel (a). (c,d) Transport scattering rate $1/\tau_{\rm tr}$ and effective mass $m^{*}$ obtained from $\sigma$ using (3); inset of Panel (c) is a scaling plot with $1/\tau_{\rm tr}$ and $\omega$ scaled by $T$, inset of Panel (d) is a scaling plotting showing $\Delta m^{*}/m=m^{*}(\omega)/m-m^{*}(0)/m$ as a function of $\omega/T$. (e,f) Modulus and phase of $\sigma(\omega)$. The effective exponent $\nu^{*}$ is explained in the text. Temperatures ranges are the same for Panels (a)-(d) and for Panels (e) and (f). Our results match the trends of the observations in Figs 3a,3b,1c,2b,1d,2d of Ref. 6.
@@
Discussion—The spatially inhomogeneous fermion-boson coupling $g^{\prime}({\bm{r}})$ in the 2d-YSYK model in (1) is an alternative to ‘random mass’ $\delta s({\bm{r}})$ disorder in Hertz-Millis theory, in the crossover regime where the bosonic modes remain extended despite the disorder [8], as we discussed below (2). The striking similarities between the results presented here and a variety of properties measured in strange metal superconductors such as the cuprates – including static and dynamic properties across the normal and superconducting states – indicate such disorder plays a significant role in the phenomenology of these systems, making it important to determine the microscopic origin of disorder configurations that strongly affect the local position of the QCP. Finally, the crucial issue of how our results evolve as the bosonic modes localize due to disorder [8] will be addressed in future work.
Acknowledgements—We thank Erez Berg, Christophe Berthod, Andrey Chubukov, Carlo Di Castro, Antoine Georges, Peter Lunts, Dmitrii Maslov, Srinivas Raghu, and Dirk van der Marel for valuable discussions. This research was supported by the U.S. National Science Foundation grant No. DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S.S.). The Flatiron Institute is a division of the Simons Foundation. J.S. and D. V. were supported by the German Research Foundation (DFG) through CRC TRR 288 “ElastoQMat,” project A07. I.E. was supported by the University of Wisconsin– Madison.